OpenStudy (cedric):

Solve -2x + 1 less than or equal to -7

OpenStudy (anonymous):

\[-2x+1 \le -7\] Subtract both sides by 1. \[-2x+1-1 \le -7-1\] \[-2x \le -8\] Divide both sides by -2. \[\frac{-2x}{-2} \le \frac{-8}{-2}\] sign of inequality changes on division or multiplication by a negative no. \[x \ge 4\]

OpenStudy (cedric):

so if this was on a graph would the circle be closed or open

OpenStudy (anonymous):

graph or number line??

OpenStudy (cedric):

number line

OpenStudy (anonymous):

shubham.bagrecha is correct

OpenStudy (anonymous):